# Improving LLM-as-a-Judge Inference with the Judgment Distribution

### v1

@@ B.2.2 Pairwise Ranking
We prompt the LLM judge to express its preference on a $K$-point Likert scale: [>,<][>,<][ > , < ] (Likert-2), [>,=,<][>,=,<][ > , = , < ] (Likert-3), or [,>,=,<,]much-greater-thanmuch-less-than[$\gg$,>,=,<$\ll$][ ≫ , > , = , < , ≪ ] (Likert-5). Assigning the symbols [,>,=,<,]much-greater-thanmuch-less-than[$\gg$,>,=,<$\ll$][ ≫ , > , = , < , ≪ ] the numerical values $[2,1,0,-1,-2]$, the methods for pairwise ranking then follow those above for pairwise scoring. We remark that the ‘mode’ and ‘mean’ for pairwise scoring and pairwise ranking in Table 1 are with post-aggregation and pre-aggregation, respectively.
@@ Appendix E Additional Results
Tables 10 ($K=9$) and 13 ($K=99$) show pointwise results over methods (expanded versions of Tables 4 and 6). Tables 11 and 12 show expanded tie analyses on RewardBench (simplified in Table 5) and MT-Bench.
Table 14 shows pairwise ranking results over methods, extending Table 7. Table 15 compares the Likert scales used for pairwise ranking, extending Table 8. In Table 15, the most calibrated setting on MT-Bench is (GPT-4o) Likert-5 no-CoT, achieving a 31% lower MSE than the most accurate setting, Likert-2 CoT, suggesting that a finer granularity has potential to improve calibration Liu et al. (2024b). With GPT-4o in Tables 14 and 15, MSE is always minimized with no-CoT, highlighting the discord between CoT’s sharpening effect and calibration. This result is in line with AlpacaEval Dubois et al. (2024), which uses no-CoT and judgment probabilities, but deviating from WB-Reward and Arena-Hard-Auto Lin et al. (2024); Li et al. (2024b), which use CoT and decoded judgments.
@@
Model Method Tie rate mean’s accuracy Non-tie accuracy $\Delta$uparrow $K=9$ $K=99$ $K=9$ $K=99$ $K=9$ $K=99$ GPT-4o mode .13, .17 .09, .20 64, 72 61, 73 +0.0, –0.1 +0.0, –0.3 [mean] .13, .16 .02, .03 65, 67 61, 53 +0.0, +0.0 +0.0, –0.0 medi .13, .17 .06, .09 65, 70 58, 62 –0.0, +0.0 –0.0, –0.1 1p .13, .16 .05, .08 66, 66 58, 60 +0.1, +0.1 +0.0, +0.5 Llama -3.1-8B mode .27, .35 .18, .24 60, 69 63, 70 +0.2, –0.2 +0.2, –2.0 [mean] .25, .26 .07, .07 58, 64 56, 61 +0.0, +0.0 +0.0, +0.0 medi .26, .29 .11, .11 60, 67 59, 67 +0.1, –0.5 –0.2, –0.6 1p .24, .23 .08, .08 61, 65 55, 57 +0.3, +0.6 +0.8, +0.1
@@
Model Method Tie rate mean’s accuracy Non-tie accuracy $\Delta$uparrow $K=9$ $K=99$ $K=9$ $K=99$ $K=9$ $K=99$ GPT-4o mode .13, .21 .08, .22 62, 62 64, 67 +0.0, –0.2 +0.1, –0.7 [mean] .13, .21 .02, .04 61, 64 42, 48 +0.0, +0.0 +0.0, +0.0 medi .13, .22 .05, .11 61, 64 64, 55 +0.0, +0.1 +0.0, –0.6 1p .14, .19 .06, .09 56, 62 67, 56 –0.1, +0.1 +0.3, +0.7 Llama -3.1-8B mode .25, .45 .14, .26 65, 71 61, 65 –0.1, –1.0 –0.4, –3.0 [mean] .24, .36 .06, .09 63, 68 49, 55 +0.0, +0.0 +0.0, –0.1 medi .25, .40 .10, .18 65, 69 54, 58 +0.0, –0.3 –0.4, +0.5 1p .20, .23 .07, .07 60, 59 53, 50 +0.1, –0.3 +1.8, +0.3
@@
GPT-4o mode 86.1${}_{\text{+1.0}}$, 81.7–${}_{\text{--2.3}}$118+.${}_{\text{+.002}}$, .134+.${}_{\text{+.016}}$ 83.8${}_{\text{+1.9}}$, 78.4–${}_{\text{--2.1}}$152+.${}_{\text{+.000}}$, .158+.${}_{\text{+.013}}$
mean 87.4+0.${}_{\text{+0.0}}$, 86.7–${}_{\text{--1.3}}$097–.${}_{\text{--.002}}$, .108+.${}_{\text{+.006}}$ 84.8${}_{\text{+1.2}}$, 82.9–${}_{\text{--0.3}}$105–.${}_{\text{--.010}}$, .099+.${}_{\text{+.002}}$
[mean] 87.0${}_{\text{+1.9}}$, 86.5${}_{\text{+1.3}}$124+.${}_{\text{+.008}}$, .127+.${}_{\text{+.018}}$ 85.1${}_{\text{+3.1}}$, 82.7${}_{\text{+2.5}}$167+.${}_{\text{+.017}}$, .182+.${}_{\text{+.036}}$
medi 86.7${}_{\text{+1.7}}$, 85.2${}_{\text{+0.6}}$119+.${}_{\text{+.003}}$, .126+.${}_{\text{+.014}}$ 84.1${}_{\text{+2.1}}$, 81.5${}_{\text{+1.3}}$160+.${}_{\text{+.010}}$, .170+.${}_{\text{+.028}}$
1p 86.6${}_{\text{+1.8}}$, 86.4${}_{\text{+2.1}}$121+.${}_{\text{+.001}}$, .116+.${}_{\text{+.000}}$ 84.2${}_{\text{+1.6}}$, 82.7${}_{\text{+1.7}}$159+.${}_{\text{+.018}}$, .165+.${}_{\text{+.027}}$
ram 87.1–${}_{\text{--0.3}}$, 86.7–${}_{\text{--1.7}}$098–.${}_{\text{--.001}}$, .104+.${}_{\text{+.004}}$ 85.1${}_{\text{+1.2}}$, 83.0–${}_{\text{--0.4}}$106–.${}_{\text{--.009}}$, .098+.${}_{\text{+.002}}$
qt 87.3–${}_{\text{--0.1}}$, 86.6–${}_{\text{--1.3}}$112+.${}_{\text{+.005}}$, .114+.${}_{\text{+.018}}$ 84.8${}_{\text{+1.3}}$, 82.7–${}_{\text{--0.5}}$149+.${}_{\text{+.010}}$, .147+.${}_{\text{+.029}}$
ps 87.3–${}_{\text{--0.1}}$, 86.6–${}_{\text{--1.2}}$105–.${}_{\text{--.001}}$, .105+.${}_{\text{+.009}}$ 84.8${}_{\text{+1.3}}$, 82.4–${}_{\text{--0.9}}$130–.${}_{\text{--.006}}$, .107+.${}_{\text{+.004}}$
Llama -3.1-8B mode 73.4${}_{\text{+3.8}}$, 72.0–${}_{\text{--0.2}}$222–.${}_{\text{--.015}}$, .221+.${}_{\text{+.029}}$ 77.3${}_{\text{+2.4}}$, 75.1${}_{\text{+3.2}}$191+.${}_{\text{+.014}}$, .169+.${}_{\text{+.027}}$
mean 75.9${}_{\text{+3.2}}$, 79.3+0.${}_{\text{+0.0}}$183–.${}_{\text{--.015}}$, .156+.${}_{\text{+.001}}$ 79.3${}_{\text{+0.6}}$, 81.3–${}_{\text{--0.2}}$125–.${}_{\text{--.004}}$, .103–.${}_{\text{--.001}}$
[mean] 75.3${}_{\text{+5.2}}$, 78.5${}_{\text{+3.5}}$229–.${}_{\text{--.009}}$, .198+.${}_{\text{+.012}}$ 79.3${}_{\text{+3.6}}$, 80.7${}_{\text{+5.7}}$201+.${}_{\text{+.029}}$, .180+.${}_{\text{+.035}}$
medi 74.4${}_{\text{+4.6}}$, 76.5${}_{\text{+2.9}}$228–.${}_{\text{--.010}}$, .207+.${}_{\text{+.016}}$ 78.4${}_{\text{+3.2}}$, 80.1${}_{\text{+6.2}}$198+.${}_{\text{+.022}}$, .161+.${}_{\text{+.019}}$
1p 76.2${}_{\text{+6.0}}$, 78.5${}_{\text{+2.5}}$218–.${}_{\text{--.020}}$, .195+.${}_{\text{+.012}}$ 80.6${}_{\text{+3.8}}$, 81.5${}_{\text{+2.3}}$187+.${}_{\text{+.015}}$, .177+.${}_{\text{+.030}}$
ram 76.1${}_{\text{+3.4}}$, 79.7–${}_{\text{--0.2}}$179–.${}_{\text{--.021}}$, .152+.${}_{\text{+.000}}$ 79.7${}_{\text{+0.9}}$, 81.1–${}_{\text{--0.3}}$123–.${}_{\text{--.007}}$, .102+.${}_{\text{+.000}}$
qt 75.7${}_{\text{+2.9}}$, 78.7–${}_{\text{--0.3}}$214–.${}_{\text{--.006}}$, .177+.${}_{\text{+.013}}$ 78.8${}_{\text{+0.1}}$, 81.3${}_{\text{+0.2}}$179+.${}_{\text{+.025}}$, .143+.${}_{\text{+.027}}$
ps 75.7${}_{\text{+2.9}}$, 78.6–${}_{\text{--0.3}}$203–.${}_{\text{--.013}}$, .163+.${}_{\text{+.002}}$ 78.6+0.${}_{\text{+0.0}}$, 81.8${}_{\text{+0.4}}$151+.${}_{\text{+.002}}$, .111+.${}_{\text{+.001}}$
@@ Appendix F Analysis
Table 17: Spearman’s $\rho$ between standard deviation of human judgments and that of LLM’s judgment distribution. Comma-separated values are with and without CoT. Bold denotes significant correlation ($\alpha=0.01$). Ranking uses Likert-3; scoring uses $K=9$ converted to a Likert-3 distribution $[P(X_{1}>X_{2}),P(X_{1}=X_{2}),P(X_{1}<X_{2})]$.
@@ F.1.3 Pluralistic Alignment
where $\Gamma(\mu,\nu)$ is the set of couplings of $\mu$ and $\nu$. A higher $p$ more heavily punishes large point distances $|x-y|$. We scale the judgment spaces to $[0,1]$ so that $W_{p}(\mu,\nu)\in[0,1]$.
@@ Results
Tables 22 and 23 report position bias in the pairwise settings. We find that no-CoT always improves MSE, even when it hurts MAE, showing that no-CoT reduces cases of extreme position bias.
@@ G.1 Approximability of Discrete Pointwise Functions Under Discretization
Let $X$ be a random variable with support $S\subset[\frac{1}{2},K+\frac{1}{2})$ for an integer $K$. Define its discretization $\hat{X}$ by $P(\hat{X}=\hat{x})\coloneq P([X]=\hat{x})$ for $\hat{x}\in\hat{S}\coloneq\{1,\ldots,K\}$, where []delimited-[][$\cdot$][ ⋅ ] denotes rounding to the nearest integer.
@@
    mode may fail to be approximated: Suppose $X$ has a density $f_{X}$ that is $L$-Lipschitz with $L\leq 1$ and achieves its supremum at $x^{*}\in\arg\max_{x\in S}f_{X}(x)$. Let $\hat{x}^{*}\in\arg\max_{\hat{x}\in\hat{S}}P(\hat{X}=\hat{x})$. Suppose $\hat{x}\in\hat{S}$, with arbitrarily large $|\hat{x}-\hat{x}^{*}|>1$, satisfies $P(\hat{X}=\hat{x}^{*})\geq P(\hat{X}=\hat{x})+\frac{L}{4}$. The above is consistent with $[x^{*}]=\hat{x}$.
@@
    If $L=0$, the claim is immediate; assume not. Define $d\coloneq\frac{L}{4}(\sqrt{1+8/L}-2)\geq\frac{L}{4}$. Let $f_{X}(x)=(d-\frac{L}{4})+L(x-\hat{x}+\frac{1}{2})$ for $x\in[\hat{x}-\frac{1}{2},\hat{x})$, and $f_{X}(x)=(d-\frac{L}{4})+L(\hat{x}-x+\frac{1}{2})$ for $x\in[\hat{x},\hat{x}+\frac{1}{2})$, and $f_{X}(x)=d+\frac{L}{4}$ for $[x]=\hat{x}^{*}$.
    Around the regions $[\hat{x}-\frac{1}{2},\hat{x}+\frac{1}{2}),[\hat{x}^{*}-\frac{1}{2},\hat{x}^{*}+\frac{1}{2})$, we let $f_{X}$ decrease to 0 with slope ±$\pm L$, or until reaching the domain boundary or each other. Continuity is maintained at the junction because, supposing $\hat{x}<\hat{x}^{*}$ without loss of generality, the nearest endpoints $\hat{x}+\frac{1}{2},\hat{x}^{*}-\frac{1}{2}$ satisfy $|(\hat{x}+\frac{1}{2})-(\hat{x}^{*}-\frac{1}{2})|\geq 1$ and $|f_{X}(\hat{x}+\frac{1}{2})-f_{X}(\hat{x}^{*}-\frac{1}{2})|=\frac{L}{2}$.
    We verify that $P(\hat{X}=\hat{x}^{*})=d+\frac{L}{4}=P(\hat{X}=\hat{x})+\frac{L}{4}$ and $\hat{x}\in\{\hat{x}\}\cup[\hat{x}^{*}-\frac{1}{2},\hat{x}^{*}+\frac{1}{2})=\arg\max_{x\in S}f_{X}(x)$.
@@
    Denote the measures of $X,\hat{X}$ as $\mu_{X},\mu_{\hat{X}}$. The definition of $(X,\hat{X})$ is equivalent to the existence of a coupling $\gamma\in\Gamma(\mu_{X},\mu_{\hat{X}})$ with samples defined by $(x,\hat{x})\sim\gamma$ for $x\sim\mu_{X}$ and $\hat{x}=[x]$.
    $|\mathbb{E}X-\mathbb{E}\hat{X}|=\left|\int(x-\hat{x})\operatorname{d}\!{\gamma}(x,\hat{x})\right|$
@@
The suppositions in (1) are to impose regularity and show even then approximation may not hold. For an example of their omission, without requiring absolutely continuous $X$, it could place atoms at arbitrary $x$, preventing any margin $P(\hat{X}=\hat{x}^{*})-P(\hat{X}=\hat{x})$ less than 1 from producing an error bound. The crucial case that causes the mode to be unstable to approximate is the case of multimodality.

### v2

@@ B.2.2 Pairwise Ranking
We prompt the LLM judge to express its preference on a $K$-point Likert scale: $[>,<]$ (Likert-2), $[>,=,<]$ (Likert-3), or [,>,=,<,][$\gg$$,\ll$] (Likert-5). Assigning the symbols [,>,=,<,][$\gg$$,\ll$] the numerical values [2,1,0,1,2][2,1,0,-1,-2], the methods for pairwise ranking then follow those above for pairwise scoring. We remark that the ‘mode’ and ‘mean’ for pairwise scoring and pairwise ranking in Table 1 are with post-aggregation and pre-aggregation, respectively.
@@ Appendix E Additional Results
Table 10 is an expanded version of Table 1, providing a subset breakdown of RewardBench. We observe particularly large gains for pointwise scoring on the Reasoning subset, e.g. absolute +7.7% and +17.1% for GPT-4o and Llama-3.1-8B.
Table 10: Mode vs. mean and CoT vs. no-CoT (comma-separated) accuracy results (%). Expanded version of Table 1.
Tables 11 ($K=9$) and 14 ($K=99$) show pointwise results over methods (expanded versions of Tables 4 and 6). Tables 12 and 13 show expanded tie analyses on RewardBench (simplified in Table 5) and MT-Bench.
Table 15 shows pairwise ranking results over methods, extending Table 7. Table 16 compares the Likert scales used for pairwise ranking, extending Table 8. In Table 16, the most calibrated setting on MT-Bench is (GPT-4o) Likert-5 no-CoT, achieving a 31% lower MSE than the most accurate setting, Likert-2 CoT, suggesting that a finer granularity has potential to improve calibration Liu et al. (2024b). With GPT-4o in Tables 15 and 16, MSE is always minimized with no-CoT, highlighting the discord between CoT’s sharpening effect and calibration. This result is in line with AlpacaEval Dubois et al. (2024), which uses no-CoT and judgment probabilities, but deviating from WB-Reward and Arena-Hard-Auto Lin et al. (2024); Li et al. (2024b), which use CoT and decoded judgments.
@@
Model Method Tie rate mean’s accuracy Non-tie accuracy $\Delta$ $\uparrow$ $K=9$ $K=99$ $K=9$ $K=99$ $K=9$ $K=99$ GPT-4o mode .13, .17 .09, .20 64, 72 61, 73 +0.0, –0.1 +0.0, –0.3 [mean] .13, .16 .02, .03 65, 67 61, 53 +0.0, +0.0 +0.0, –0.0 medi .13, .17 .06, .09 65, 70 58, 62 –0.0, +0.0 –0.0, –0.1 1p .13, .16 .05, .08 66, 66 58, 60 +0.1, +0.1 +0.0, +0.5 Llama -3.1-8B mode .27, .35 .18, .24 60, 69 63, 70 +0.2, –0.2 +0.2, –2.0 [mean] .25, .26 .07, .07 58, 64 56, 61 +0.0, +0.0 +0.0, +0.0 medi .26, .29 .11, .11 60, 67 59, 67 +0.1, –0.5 –0.2, –0.6 1p .24, .23 .08, .08 61, 65 55, 57 +0.3, +0.6 +0.8, +0.1
@@
Model Method Tie rate mean’s accuracy Non-tie accuracy $\Delta$ $\uparrow$ $K=9$ $K=99$ $K=9$ $K=99$ $K=9$ $K=99$ GPT-4o mode .13, .21 .08, .22 62, 62 64, 67 +0.0, –0.2 +0.1, –0.7 [mean] .13, .21 .02, .04 61, 64 42, 48 +0.0, +0.0 +0.0, +0.0 medi .13, .22 .05, .11 61, 64 64, 55 +0.0, +0.1 +0.0, –0.6 1p .14, .19 .06, .09 56, 62 67, 56 –0.1, +0.1 +0.3, +0.7 Llama -3.1-8B mode .25, .45 .14, .26 65, 71 61, 65 –0.1, –1.0 –0.4, –3.0 [mean] .24, .36 .06, .09 63, 68 49, 55 +0.0, +0.0 +0.0, –0.1 medi .25, .40 .10, .18 65, 69 54, 58 +0.0, –0.3 –0.4, +0.5 1p .20, .23 .07, .07 60, 59 53, 50 +0.1, –0.3 +1.8, +0.3
@@
[…]
@@
E.1 DeepSeek-V3 Results
We provide partial results for DeepSeek-V3 DeepSeek-AI et al. (2025), a model of comparable size to GPT-4o. Tables 17 and 18 contain pointwise and listwise results, respectively. The trends for DeepSeek-V3 match those of GPT-4o.
Method Acc $\uparrow$ MSE $\downarrow$
mode 84.7, 82.5 0.123, 0.128
mean 84.8, 84.2 0.119, 0.120
[mean] 84.7, 82.7 0.123, 0.127
medi 84.7, 82.6 0.123, 0.128
1p 84.5, 82.9 0.124, 0.126
ram 84.9, 84.1 0.120, 0.118
qt 85.0, 83.9 0.122, 0.125
ps 85.0, 83.9 0.122, 0.125
Table 17: Pointwise results with DeepSeek-V3 on RewardBench. $K=9$. Comma-separated values are with and without CoT. Text styling follows Table 1.
Space Method Nectar RM-Bench
Acc MSE Acc MSE
direct list mode 83.6 0.149 67.8 0.322
direct list mean 84.0 0.129 67.6 0.307
Table 18: Listwise results with DeepSeek-V3. Text styling follows Table 1.
@@ Appendix F Analysis
Table 20: Spearman’s $\rho$ between standard deviation of human judgments and that of LLM’s judgment distribution. Comma-separated values are with and without CoT. Bold denotes significant correlation ($\alpha=0.01$). Ranking uses Likert-3; scoring uses $K=9$ converted to a Likert-3 distribution $P(X_{1}>X_{2}),P(X_{1}=X_{2}),P(X_{1}<X_{2})]$.
@@ F.1.3 Pluralistic Alignment
where $\Gamma(\mu,\nu)$ is the set of couplings of $\mu$ and $\nu$. A higher $p$ more heavily punishes large point distances |xy||x-y|. We scale the judgment spaces to $[0,1]$ so that $W_{p}(\mu,\nu)\in[0,1]$.
@@ Results
Tables 25 and 26 report position bias in the pairwise settings. We find that no-CoT always improves MSE, even when it hurts MAE, showing that no-CoT reduces cases of extreme position bias.
@@ G.1 Approximability of Discrete Pointwise Functions Under Discretization
[…]
@@
[…]
@@
[…]
    Around the regions $\hat{x}-\frac{1}{2},\hat{x}+\frac{1}{2}),[\hat{x}^{*}-\frac{1}{2},\hat{x}^{*}+\frac{1}{2})$, we let $f_{X}$ decrease to 0 with slope $\pm L$, or until reaching the domain boundary or each other. Continuity is maintained at the junction because, supposing $\hat{x}<\hat{x}^{*}$ without loss of generality, the nearest endpoints $\hat{x}+\frac{1}{2},\hat{x}^{*}-\frac{1}{2}$ satisfy $|(\hat{x}+\frac{1}{2})-(\hat{x}^{*}-\frac{1}{2})|\geq 1$ and $|f_{X}(\hat{x}+\frac{1}{2})-f_{X}(\hat{x}^{*}-\frac{1}{2})|=\frac{L}{2}$.
[…]
@@
[…]
@@
[…]